OpenStudy (anonymous):

If sin( sin inverse 1/5 + cos inverse x) = 1 Find the value of x

OpenStudy (rational):

you may use sin(A+B) formula

OpenStudy (anonymous):

That is sin(A + B) = sinAcosB + SinBcosA?

OpenStudy (anonymous):

@rational Uhm if SinA is sin^-1 1/5 what is SinB?

OpenStudy (rizags):

i'd focus on the fact that sin(....)=1, specifically the fact that it equals 1

OpenStudy (anonymous):

So we get it into a sin^2 x + cos^2 x form?

OpenStudy (rational):

\(\sin(\sin^{-1}x) = x\)

OpenStudy (rizags):

look at it this way: \[if \sin^{-1} \frac{ 1 }{ 5 }\rightarrow \sin y=\frac{ 1 }{ 5 }, and \cos^{-1} x \rightarrow \cos z=x\]

OpenStudy (anonymous):

Oh okay so if sin y = 1/5 cos z = x I write it as Sin(sin y 1/5 . cos z + sin z 1/5 . cos x) = 1?

OpenStudy (dan815):

oh nice trick

OpenStudy (dan815):

use the sin(A+B)

OpenStudy (anonymous):

sec writing it in my book

OpenStudy (rational):

Alternatively if you can recall some inverse identities, start like this\[\sin^{-1}(1/5) + \cos^{-1}(x) = \dfrac{\pi}{2}\]

OpenStudy (rational):

(assuming principle angles.. )

OpenStudy (rational):

Notice that x=1/5 satisfies above equation if you remember the identity \[\sin^{-1}x + \cos^{-1}x = \dfrac{\pi}{2}\]

OpenStudy (anonymous):

I got this so far.. sin(sin^-1 1/5).cos(cos^-1 x) + sin(cos^-1 x).cos(sin^-1 1/5) = 1 And that = 1/5 . x + sin(cos^-1 x).cos(sin^-1 1/5)

OpenStudy (anonymous):

Yeah I'm stuck. How would you use the pi/2 when cos inverse and sin inverse are locked in the round brackets?

OpenStudy (rational):

fine so far, next eliminate trig completely.. cos(sin^-1 1/5) = sqrt(24)/5

OpenStudy (rational):

sin(cos^-1 x) = sqrt(1-x^2)

OpenStudy (anonymous):

How'd you get sqrt(24)/5?

OpenStudy (rational):

\(\sin^{-1}(1/5) = t \implies \sin t = 1/5 \implies \cos t = \sqrt{1-(1/5)^2} = \sqrt{24}/5\)

OpenStudy (anonymous):

Oh.. oooh o. Okay I get it! hang on, lemme try completing the rest of the problem

OpenStudy (rational):

Okay... rest is just algebra

OpenStudy (ybarrap):

$$ \sin\left ( \sin^{-1}{1/5 }+ \cos ^{-1} x\right ) = 1\\ \sin^{-1}\sin(\left ( \sin^{-1}{1/5 }+ \cos ^{}-1 x\right ))=\sin^{-1}1={\pi\over2}\\ \cos ^{-1} x={\pi\over2}-\sin^{-1}{1\over 5 }\\ x=\cos\left({\pi\over2}-\sin^{-1}{1\over 5 }\right)\\ $$ Does this help?

OpenStudy (anonymous):

Got it.. finally its 1/5 the inverse cancels sin and cos x/5 + root1 + x^2.root24/5 = 1 multiplying both sides by 5 and subtracting x we get root 24/5.root 1 - x^2 = 5 - x squaring both sides 2.6.1-x^2 = 25 - x^2 - 25 - 10x or 24 - 24x^2 = 25 + x^2 - 10x or 25x^2 + 1 + 10x = 0 (5x -1)^2 = 0 5x - 1 = 0 x = 1/5

OpenStudy (anonymous):

Thanks everyone :) @ybarrap @rational @dan815

OpenStudy (rational):

Wow! you managed to sovle that equation congrats!! look at ybarrap's reply, it is the shortest and most straightforward solution